# JWST PRIMER: A lack of outshining in four normal galaxies from the ALMA-CRISTAL Survey

### v1

@@ 5.2 The rest-UV and optical sizes
Another way to probe the morphology of old and young stellar light is to compare the observed sizes of galaxies above and below the Balmer break. Ono etĀ al. (2024) andĀ Morishita etĀ al. (2024) showed that the ratio of the rest-optical to UV size was approximately unity in a sample of galaxies from ā and ā respectively. Ono etĀ al. (2024) take this observation as evidence that these galaxies are being caught at the start of inside-out growth, where the young and old stars have yet to be spatially differentiated. A key complication in this comparison comes from the complex and diverse morphologies of the CRISTAL galaxies (and high-redshift galaxies in general). Indeed, in the analysis ofĀ Varadaraj etĀ al. (2024), all four of the galaxies we study here were present in the parent sample but failed in the SĆ©rsic fitting step due to poor residuals. It is apparent in our analysis that the CRISTAL galaxies cannot be adequately described with a single SĆ©rsic model, and hence we quote the non-parametric sizes fromĀ Varadaraj etĀ al. (2024) instead. As shown in TableĀ 3, we find that the rest-frame UV and optical sizes are consistent within the errors, with the exception of CRISTAL-13. CRISTAL-13 has the most extended and complex morphology of our sample, and also potentially harbours a disk like feature in the WT (although this is blue/young and does not represent the centre of mass of the system). The sizes of the CRISTAL sources are typical for their mass, with half-light radii of around . Hence we see no evidence for different rest-frame UV and optical sizes, that might indicate inside-out growth being established or multiple stellar population ages that are significantly spatially decoupled. As we have previously discussed, the half-light radius does not fully describe the extent of the galaxy, which is of the order of ā at as we have shown in Fig.Ā 1. The size comparison here further supports the visually identified similar morphology across the broad observed wavelength range probed byĀ JWST NIRCam. We discuss further in SectionĀ 5.4 the differences in size and morphology between the rest-frame UV/optical light and the observed FIR dust continuum emission.
@@ 5.3 A lack of outshining in the CRISTAL galaxies
With our resolved SED fitting analysis using the PRIMER data, we can test the impact of outshining in galaxies of higher stellar mass, which arguably could be expected to show a more established stellar population and hence a smaller impact of extremely young bursts that leads to outshining. We compare the stellar masses from the resolved and integrated fits in Fig.Ā 5. Here we show the results from our fiducial analysis, which assumed a CSFH, and the results from assuming a delayed- SFH for comparison. The results are consistent between these two different SFHs, providing reassurance in our results and allowing a more direct comparison to previous studies (who typically used one of these parameterizations). We compare to the five galaxies studied byĀ GimĆ©nez-Arteaga etĀ al. (2023), which were gravitationally lensed (we show the delensed masses), and to the sample of star-forming galaxies at ā byĀ Shen etĀ al. (2024). Shen etĀ al. (2024) analyse galaxies with using a resolved and integrated approach, assuming a delayed exponential SFH. We also compare to the red galaxies studied byĀ PĆ©rez-GonzĆ”lez etĀ al. (2023) at ā who assumed a delayed- SFH. We find that in the three CRISTAL galaxies we study with we recover a higher mass in the resolved SED fitting analysis, but only by ā. For CRISTAL-17, the lowest mass (and highest redshift) source in our sample, we see a larger deviation. Hence we are able to bridge between the studies ofĀ GimĆ©nez-Arteaga etĀ al. (2023) at and the higher mass studies ofĀ Shen etĀ al. (2024) andĀ PĆ©rez-GonzĆ”lez etĀ al. (2023), showing the onset of outshining at the lower mass end. Our conclusions are unchanged if we instead assume a delayed- SFH, with the recovered stellar masses being within the errors of the CSFH fiducial case. Although in general the delayed- models produce closer resolved and integrated Ā estimates.
@@
In comparison to Giménez-Arteaga et al. (2023), we find that the mass-weighted stellar ages of our galaxies in the integrated analysis are significantly older (50-130 Myrs vs. the in that study) and crucially the derived integrated ages are not dramatically different from the average resolved mass-weighted age. This results in comparable SFRs and  values between the resolved and integrated fits for the CRISTAL galaxies. As shown in Fig. 5 our results, coupled with previous studies, show that at the resolved and integrated stellar masses are in good agreement, whereas at lower masses the effect of outshining can be seen and the two values diverge (e.g. for CRISTAL-17). We find lower rest-frame optical emission line equivalent-widths than the Giménez-Arteaga et al. (2023) galaxies, which is directly related to the older ages (than the lower-mass lensed sources studied in Giménez-Arteaga et al., 2023, 2024; Fujimoto et al., 2024) we recover in general for the CRISTAL sources. Even for individual clumps, we do not find any regions that exceed (H + [OIII]) à  (with the possible exception of CRISTAL-15), however Giménez-Arteaga et al. (2023) found regions with (H + [OIII])à . Future spectroscopic observations of the CRISTAL sample with the JWST NIRSpec IFU will provide resolved emission line strengths and thus a more detailed analysis of the impact of these lines on the photometry and refined estimates of the stellar ages.
@@ 5.4 The resolved IRX- relation
The presence of dust in the CRISTAL galaxies can be ascertained in two ways, first through the observation of reddening of the rest-frame UV light and second with the detection of emission directly from the dust in the FIR. At low redshift the energy balance between the Ā emanating from the dust with the absorbed Ā from young stars has been shown to lead to a correlation between the ratio of the FIR luminosity to the UV luminosity (infrared-excess /]) and its rest-frame UV slope (e.g. the seminal work byĀ Meurer etĀ al., 1999; Calzetti etĀ al., 2000). Targeted observations such as the ALMA-ALPINE and ALMA-REBELS large programs have obtained measurements of the infrared luminosity for samples of high-redshift galaxies, enabling the investigation of the evolution of the ā relation with properties such as redshift and stellar mass (e.g. Fudamoto etĀ al., 2020a, b; Bowler etĀ al., 2018, 2024). With the improved spatial resolution now available in both the rest-frame UV and FIR withĀ JWST and ALMA, it is now possible to investigate if this relation holds on spatially resolved scales and understand how the distribution of dust within galaxies affects global measurements. This may help explain the low values of galaxies such as those at studied byĀ Capak etĀ al. (2015), who concluded that the dust properties were evolving with redshift (e.g. to a lower-metallicity, Small Magellanic Cloud - SMC-like dust). Spatially resolved analysis can potentially reveal whether the position of these galaxies on the ā diagram is due to dust temperature increasing with redshift or geometric offsets between the stars and dust within the galaxyĀ (Barisic etĀ al., 2017; Faisst etĀ al., 2017; Bowler etĀ al., 2022).
@@
With the small sample of galaxies we study, that have bothĀ JWST PRIMER and ALMA-CRISTAL high-resolution observations, we can make some of the first measurements of the resolved āĀ measurement within galaxies at . CRISTAL-11 and CRISTAL-13 are detected in the rest-frame 158m dust continuumĀ (Mitsuhashi etĀ al., 2023). These detections are shown by the contours on the map in Fig.Ā 2. The results of our resolved ā analysis using the apertures at the positions shown in Fig.Ā 6 for these two galaxies are shown in Fig.Ā 7. The apertures were chosen to enclose the flux from the distinct regions. We measured a dust-associated ā value using a circular aperture centred on the peak of the ALMA dust continuum detection of diameters 0.30 arcsec for CRISTAL-11 and 0.48 arcsec for CRISTAL-13. A second value was obtained from an aperture encompassing the remaining emission. For each aperture, the resolved measurement was obtained by taking the mean of the values from the map shown in Fig.Ā 2. The error bars correspond to the range of values in the aperture, and thus represent the spread in resolved . The fraction of the rest-frame UV flux enclosed by the aperture was calculated from the F090W image and was used to scale the to obtain the aperture rest-frame UV luminosity. For the aperture centred on the ALMA dust emission peak, is taken from Mitsuhashi etĀ al. (2023) who exploit the same approach as used in the ALPINE analysisĀ (BĆ©thermin etĀ al., 2020) using a SED template derived fromĀ BĆ©thermin etĀ al. (2017) (equivalent to a modified blackbody with , ) integrated over ā. For the additional aperture, CRISTAL-11 has a detection in the dust continuum map on the south-east side of the galaxy, so this is presented as a detection in Fig.Ā 7. CRISTAL-13 has no detection associated with the western component (the ātailā), so this is presented as a (conservative) upper limit.
@@
We find that the rest-frame UV slope of all four sources are bluer than the previous values computed inĀ Faisst etĀ al. (2020) as shown in TableĀ 2. TheĀ Faisst etĀ al. (2020) values were derived using the same method we exploit, fitting a power law to the best-fitting SED model to the photometry. Thus, we find that the global IRX values are shifted bluewards in the āĀ plane, now showing good agreement with theĀ Calzetti etĀ al. (2000) starburst relation. The resolved measurements also appear to follow the predictions of this starburst dust law, although they do lie slightly above this relation potentially potentially due to further geometric effects on the scale of these clumps (e.g. seeĀ Popping etĀ al., 2017). These results are in agreement with the statistical measure of the āĀ relations presented inĀ Bowler etĀ al. (2024), where they found a Calzetti-like curve provided a good description of the REBELS and ALPINE samples. A major caveat in this analysis is the uncertain dust SED at high redshift, which as discussed inĀ Bowler etĀ al. (2024) andĀ Mitsuhashi etĀ al. (2023), can lead to systematic offsets in IRX of up to 0.4 dex with a 10K difference in the assumed (keeping the emissivity index, , fixed). Individual dust temperature measurements are required to understand the exact āĀ relation, for example recent work on HZ10 byĀ Villanueva etĀ al. (2024), however our results demonstrate that an āĀ relationship similar to the local starburst relation ofĀ Calzetti etĀ al. (2000) appears to hold at a resolved level at .
@@ 5.5 The role of dust in shaping the rest-UV morphology
Given the excellent spatial resolution of the ALMA-CRISTAL data, we can make further tentative investigations into the location of the dust within CRISTAL-11 and CRISTAL-13 in comparison to the rest-frame UV light. In Fig.Ā 6 we show the CRISTAL contours in comparison to the bluest (and hence highest resolution)Ā JWST NIRCam band, F090W. This comparison shows that the CRISTAL-11 peak dust detection lies offset by approximately 0.6 arcsec from the rest-frame UV peak in this source, the same magnitude of offset that is found for 30 percent of the ALPINE sample in the statistical analysis ofĀ Killi etĀ al. (2024). We confirm that this rest-frame UV to FIR spatial offset is due to a young star-forming clump away from the stellar mass centroid, which is also the dustiest region of the galaxy (as revealed by the SED fitting derived Ā and the direct detection of the dust continuum).
We find that in CRISTAL-11 the offset between the rest-FIR and UV emission can be naturally explained by differences in amount of dust attenuation (there is also a difference in age between these two clumps but it is small, only ). In CRISTAL-13, the dust continuum detection is offset from the bluest regions of the galaxy (in the Western tail), and instead follows the Ā map closely. If we closely inspect the map of CRISTAL-13 shown in Fig.Ā 6, we find a small offset between the ALMA dust continuum detection for CRISTAL-13 and the rest-frame UV bright NE clump of the order of 0.15 arcsec. If the dust continuum detection is in fact originating from the region offset from the NE clump, then this could point to dust attenuation leading to the gap in the observed rest-frame UV emission from this source. This dust-star morphology is similar to that seen in the āCosmic Grapesā sources, whereĀ GimĆ©nez-Arteaga etĀ al. (2024) find that the observed clumps appear to have similar ages but differ in their derived , indicating that differing dust coverage does shape the observed rest-frame UV morphology (although seeĀ Fujimoto etĀ al. 2024; Zanella etĀ al. 2024 who suggest a reduced importance of dust). These results are in close agreement with the recent simulation work ofĀ Ocvirk etĀ al. (2024) andĀ Nakazato etĀ al. (2024) (see SectionĀ 6), who found that dust attenuation can shape the observed rest-frame UV morphology, and lead to offsets between the UV and FIR peak.
@@
NE | -19.92 | ā | |||||||
WT | -20.94 | ā | |||||||
CRISTAL-15 | Global | -21.90 | ā | ||||||
CRISTAL-17 | Global | -21.37 | ā |
Given the beam size and SNR of the ALMA detection for the CRISTAL-13 source, the error on the position of the peak can be approximated as . FromĀ Mitsuhashi etĀ al. (2023) this ALMA detection is compact, with a measured size of arcsec. Therefore, we can use a smaller aperture of diameter 0.12 arcsec to determine the Ā and , centred on the peak of the ALMA detection. In this smaller apertures, we unsurprisingly recover a redder rest-frame UV slope and a higher IRX as shown with the open circle in Fig.Ā 7. This point lies above the canonical āĀ relations, indicating a highly obscured region where the observed rest-frame UV colour is not adequately describing the obscured SFR fraction, for example in the case of a dust screen with holes. A similar offset is also found for regions of the luminous galaxy HZ10 as presented inĀ Villanueva etĀ al. (2024), where they are able to derive resolved dust temperature estimates thus reducing the uncertainty on the IRX. We caution that the resolution of the ALMA data is insufficient to reliably ascertain if these features are spatially offset, however this analysis demonstrates that an inhomogeneous dust distribution within CRISTAL-11 and CRISTAL-13 can explain the observed rest-frame UV morphology, with differences in age being a secondary effect. This hypothesis is also supported by the observed negative correlation between the rest-frame UV to FIR offsets and the obscured fraction of star formation found for the full CRISTAL sampleĀ (Mitsuhashi etĀ al., 2023).

### v2

@@ 5.2 The rest-UV and optical sizes
Another way to probe the morphology of old and young stellar light is to compare the observed sizes of galaxies above and below the Balmer break. Ono et al. (2024) and Morishita et al. (2024) showed that the ratio of the rest-optical to UV size was approximately unity in a sample of galaxies from – and – respectively. Ono et al. (2024) take this observation as evidence that these galaxies are being caught at the start of inside-out growth, where the young and old stars have yet to be spatially differentiated. A key complication in this comparison comes from the complex and diverse morphologies of the CRISTAL galaxies (and high-redshift galaxies in general). Indeed, in the analysis of Varadaraj et al. (2024), all four of the galaxies we study here were present in the parent sample but failed in the Sérsic fitting step due to poor residuals. It is apparent in our analysis that the CRISTAL galaxies cannot be adequately described with a single Sérsic model, and hence we quote the non-parametric sizes from Varadaraj et al. (2024) instead. As shown in Table 3, we find that the rest-frame UV and optical sizes are consistent within the errors, with the exception of CRISTAL-13. CRISTAL-13 has the most extended and complex morphology of our sample, and also potentially harbours a disk like feature in the WT (although this is blue/young and does not represent the centre of mass of the system). The sizes of the CRISTAL sources are typical for their mass, with half-light radii of around . Hence we see no evidence for different rest-frame UV and optical sizes, that might indicate inside-out growth being established or multiple stellar population ages that are significantly spatially decoupled. As we have previously discussed, the half-light radius does not fully describe the extent of the galaxy, which is of the order of – at as we have shown in Fig. 1. The size comparison here further supports the visually identified similar morphology across the broad observed wavelength range probed by JWST NIRCam. We discuss further in Section 5.4 the differences in size and morphology between the rest-frame UV/optical light and the observed FIR dust continuum emission.
@@ 5.3 A lack of outshining in the CRISTAL galaxies
With our resolved SED fitting analysis using the PRIMER data, we can test the impact of outshining in galaxies of higher stellar mass, which arguably could be expected to show a more established stellar population and hence a smaller impact of extremely young bursts that leads to outshining. We compare the stellar masses from the resolved and integrated fits in Fig. 5. Here we show the results from our fiducial analysis, which assumed a CSFH, and the results from assuming a delayed- SFH for comparison. The results are consistent between these two different SFHs, providing reassurance in our results and allowing a more direct comparison to previous studies (who typically used one of these parameterizations). We compare to the five galaxies studied by Giménez-Arteaga et al. (2023), which were gravitationally lensed (we show the delensed masses), and to the sample of star-forming galaxies at – by Shen et al. (2024). Shen et al. (2024) analyse galaxies with using a resolved and integrated approach, assuming a delayed exponential SFH. We also compare to the red galaxies studied by Pérez-González et al. (2023) at – who assumed a delayed- SFH. We find that in the three CRISTAL galaxies we study with we recover a higher mass in the resolved SED fitting analysis, but only by –. For CRISTAL-17, the lowest mass (and highest redshift) source in our sample, we see a larger deviation. Hence we are able to bridge between the studies of Giménez-Arteaga et al. (2023) at and the higher mass studies of Shen et al. (2024) and Pérez-González et al. (2023), showing the onset of outshining at the lower mass end. Our conclusions are unchanged if we instead assume a delayed- SFH, with the recovered stellar masses being within the errors of the CSFH fiducial case. Although in general the delayed- models produce closer resolved and integrated estimates.
@@
In comparison to Giménez-Arteaga et al. (2023), we find that the mass-weighted stellar ages of our galaxies in the integrated analysis are significantly older (50-130 Myrs vs. the in that study) and crucially the derived integrated ages are not dramatically different from the average resolved mass-weighted age. This results in comparable SFRs and values between the resolved and integrated fits for the CRISTAL galaxies. As shown in Fig. 5 our results, coupled with previous studies, show that at the resolved and integrated stellar masses are in good agreement, whereas at lower masses the effect of outshining can be seen and the two values diverge (e.g. for CRISTAL-17). We find lower rest-frame optical emission line equivalent-widths than the Giménez-Arteaga et al. (2023) galaxies, which is directly related to the older ages (than the lower-mass lensed sources studied in Giménez-Arteaga et al., 2023, 2024; Fujimoto et al., 2024) we recover in general for the CRISTAL sources. Even for individual clumps, we do not find any regions that exceed (H + [OIII]) Å (with the possible exception of CRISTAL-15), however Giménez-Arteaga et al. (2023) found regions with (H + [OIII])Å. Future spectroscopic observations of the CRISTAL sample with the JWST NIRSpec IFU will provide resolved emission line strengths and thus a more detailed analysis of the impact of these lines on the photometry and refined estimates of the stellar ages.
@@ 5.4 The resolved IRX- relation
The presence of dust in the CRISTAL galaxies can be ascertained in two ways, first through the observation of reddening of the rest-frame UV light and second with the detection of emission directly from the dust in the FIR. At low redshift the energy balance between the emanating from the dust with the absorbed from young stars has been shown to lead to a correlation between the ratio of the FIR luminosity to the UV luminosity (infrared-excess /]) and its rest-frame UV slope (e.g. the seminal work by Meurer et al., 1999; Calzetti et al., 2000). Targeted observations such as the ALMA-ALPINE and ALMA-REBELS large programs have obtained measurements of the infrared luminosity for samples of high-redshift galaxies, enabling the investigation of the evolution of the – relation with properties such as redshift and stellar mass (e.g. Fudamoto et al., 2020a, b; Bowler et al., 2018, 2024). With the improved spatial resolution now available in both the rest-frame UV and FIR with JWST and ALMA, it is now possible to investigate if this relation holds on spatially resolved scales and understand how the distribution of dust within galaxies affects global measurements. This may help explain the low values of galaxies such as those at studied by Capak et al. (2015), who concluded that the dust properties were evolving with redshift (e.g. to a lower-metallicity, Small Magellanic Cloud - SMC-like dust). Spatially resolved analysis can potentially reveal whether the position of these galaxies on the – diagram is due to dust temperature increasing with redshift or geometric offsets between the stars and dust within the galaxy (Barisic et al., 2017; Faisst et al., 2017; Bowler et al., 2022).
@@
With the small sample of galaxies we study, that have both JWST PRIMER and ALMA-CRISTAL high-resolution observations, we can make some of the first measurements of the resolved – measurement within galaxies at . CRISTAL-11 and CRISTAL-13 are detected in the rest-frame 158m dust continuum (Mitsuhashi et al., 2023). These detections are shown by the contours on the map in Fig. 2. The results of our resolved – analysis using the apertures at the positions shown in Fig. 6 for these two galaxies are shown in Fig. 7. The apertures were chosen to enclose the flux from the distinct regions. We measured a dust-associated – value using a circular aperture centred on the peak of the ALMA dust continuum detection of diameters 0.30 arcsec for CRISTAL-11 and 0.48 arcsec for CRISTAL-13. A second value was obtained from an aperture encompassing the remaining emission. For each aperture, the resolved measurement was obtained by taking the mean of the values from the map shown in Fig. 2. The error bars correspond to the range of values in the aperture, and thus represent the spread in resolved . The fraction of the rest-frame UV flux enclosed by the aperture was calculated from the F090W image and was used to scale the to obtain the aperture rest-frame UV luminosity. For the aperture centred on the ALMA dust emission peak, is taken from Mitsuhashi et al. (2023) who exploit the same approach as used in the ALPINE analysis (Béthermin et al., 2020) using a SED template derived from Béthermin et al. (2017) (equivalent to a modified blackbody with , ) integrated over –. The dust continuum sizes measured in Mitsuhashi et al. (2023) for CRISTAL-11 and CRISTAL-13 are arcsec and arcsec, respectively. Comparing these to the ALMA beam sizes we see that the dust continuua are barely resolved, suggesting the emission is coming from a compact region, smaller than the aperture sizes. For the additional aperture, CRISTAL-11 has a detection in the dust continuum map on the south-east side of the galaxy, so this is presented as a detection in Fig. 7. CRISTAL-13 has no detection associated with the western component (the ‘tail’), so this is presented as a (conservative) upper limit.
@@
We find that the rest-frame UV slope of all four sources are bluer than the previous values computed in Faisst et al. (2020) as shown in Table 2. The Faisst et al. (2020) values were derived using the same method we exploit, fitting a power law to the best-fitting SED model to the photometry. Thus, we find that the global IRX values are shifted bluewards in the – plane, now showing good agreement with the Calzetti et al. (2000) starburst relation. The resolved measurements also appear to follow the predictions of this starburst dust law, although they do lie slightly above this relation potentially potentially due to further geometric effects on the scale of these clumps (e.g. see Popping et al., 2017). These results are in agreement with the statistical measure of the – relations presented in Bowler et al. (2024), where they found a Calzetti-like curve provided a good description of the REBELS and ALPINE samples. One caveat in using the – relation is the unknown intrinsic rest-frame UV slope of the sample, which affects the position of the intercept of the relation with the -axis. Thus, we show the fitted relations from Bowler et al. (2024) for both and , since these values denote the range of values inferred from the SED fits to the REBELS and ALPINE sources. Another major caveat in this analysis is the uncertain dust SED at high redshift, which as discussed in Bowler et al. (2024) and Mitsuhashi et al. (2023), can lead to systematic offsets in IRX of up to 0.4 dex with a 10K difference in the assumed (keeping the emissivity index, , fixed). Individual dust temperature measurements are required to understand the exact – relation, for example recent work on HZ10 by Villanueva et al. (2024), however our results demonstrate that an – relationship similar to the local starburst relation of Calzetti et al. (2000) appears to hold at a resolved level at .
@@ 5.5 The role of dust in shaping the rest-UV morphology
Given the excellent spatial resolution of the ALMA-CRISTAL data, we can make further tentative investigations into the location of the dust within CRISTAL-11 and CRISTAL-13 in comparison to the rest-frame UV light. The other two sources we study were not detected in the ALMA-CRISTAL survey in the dust continuum (Mitsuhashi et al., 2023). From the ALPINE and REBELS surveys it has been shown that the dust obscured SFR (or obscured fraction) drops rapidly around (Fudamoto et al., 2020b; Algera et al., 2023; Bowler et al., 2024), and hence the fact that CRISTAL-15 and CRISTAL-17 are undetected (especially given the low S/N of the other detections) is unsurprising. Furthermore, these sources are bluer than CRISTAL-11 and CRISTAL-13, with CRISTAL-17 also having a significantly lower SFR and higher redshift. Despite this, we see regions of comparable across the sources, and deeper ALMA data may yet reveal dust continuum detections at these positions.
In Fig. 6 we show the CRISTAL contours in comparison to the bluest (and hence highest resolution) JWST NIRCam band, F090W. This comparison shows that the CRISTAL-11 peak dust detection lies offset by approximately 0.6 arcsec from the rest-frame UV peak in this source, the same magnitude of offset that is found for 30 percent of the ALPINE sample in the statistical analysis of Killi et al. (2024). We confirm that this rest-frame UV to FIR spatial offset is due to a young star-forming clump away from the stellar mass centroid, which is also the dustiest region of the galaxy (as revealed by the SED fitting derived and the direct detection of the dust continuum).
We find that in CRISTAL-11 the offset between the rest-FIR and UV emission can be naturally explained by differences in amount of dust attenuation (there is also a difference in age between these two clumps but it is small, only ). In CRISTAL-13, the dust continuum detection is offset from the bluest regions of the galaxy (in the Western tail), and instead follows the map closely. If we closely inspect the map of CRISTAL-13 shown in Fig. 6, we find a small offset between the ALMA dust continuum detection for CRISTAL-13 and the rest-frame UV bright NE clump of the order of 0.15 arcsec. If the dust continuum detection is in fact originating from the region offset from the NE clump, then this could point to dust attenuation leading to the gap in the observed rest-frame UV emission from this source. This dust-star morphology is similar to that seen in the ‘Cosmic Grapes’ sources, where Giménez-Arteaga et al. (2024) find that the observed clumps appear to have similar ages but differ in their derived , indicating that differing dust coverage does shape the observed rest-frame UV morphology (although see Fujimoto et al. 2024; Zanella et al. 2024 who suggest a reduced importance of dust). These results are in close agreement with the recent simulation work of Ocvirk et al. (2024) and Nakazato et al. (2024) (see Section 6), who found that dust attenuation can shape the observed rest-frame UV morphology, and lead to offsets between the UV and FIR peak.
@@
NE | -19.92 | – | |||||||
WT | -20.94 | – | |||||||
CRISTAL-15 | Global | -21.90 | – | ||||||
CRISTAL-17 | Global | -21.37 | – |
Given the beam size and SNR of the ALMA detection for the CRISTAL-13 source, the error on the position of the peak can be approximated as . From Mitsuhashi et al. (2023) this ALMA detection is compact, with a measured size of arcsec. Therefore, we can use a smaller aperture of diameter 0.12 arcsec to determine the and , centred on the peak of the ALMA detection. We correct for the effect of aperture loss in the F090W band due to the PSF using a factor of . In this smaller aperture, we unsurprisingly recover a redder rest-frame UV slope and a higher IRX as shown with the open circle in Fig. 7. This point lies above the canonical – relations, indicating a highly obscured region where the observed rest-frame UV colour is not adequately describing the obscured SFR fraction, for example in the case of a dust screen with holes. A similar offset is also found for regions of the luminous galaxy HZ10 as presented in Villanueva et al. (2024), where they are able to derive resolved dust temperature estimates thus reducing the uncertainty on the IRX. We caution that the resolution of the ALMA data is insufficient to reliably ascertain if these features are spatially offset, however this analysis demonstrates that an inhomogeneous dust distribution within CRISTAL-11 and CRISTAL-13 can explain the observed rest-frame UV morphology, with differences in age being a secondary effect. This hypothesis is also supported by the observed negative correlation between the rest-frame UV to FIR offsets and the obscured fraction of star formation found for the full CRISTAL sample (Mitsuhashi et al., 2023).